Investment Theory:

Assume corporate taxes as detailed in the following question:

6. An all-equity firm has 155,000 shares of common stock outstanding, currently worth $20 per share. Its equity holders require a 20% return. The firm decides to issue $1 million of 10% debt and use the proceeds to repurchase common stock. The corporate tax rate is 30%.

a. What is the market value of the firm before the repurchase?

b. According to Modigliani-Miller, what is the market value of the firm after the repurchase?

c. What is the market value of the firm’s equity after the repurchase?
Answer #1

a. Market Value of firm =Number of shares *Price of share =155000*20 =3,110,000
b. Market Value of the firm after repurchase =Market value of all equity firm+Debt*Tax rate =3110000+1000000*30% =3,410,000
c. Market Value of firm's equity after repurchase =Market Value of the firm after repurchase-debt =3410000-1000000 =2,410,000